OpenStudy (anonymous):

Determine the number of real solutions of -2x^2 + 5x - 3 = 0. A. 0 B. 1 (double root) C. 2

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

what did you get

jimthompson5910 (jim_thompson5910):

you can either factor or you can use the quadratic formula

OpenStudy (anonymous):

I got B

jimthompson5910 (jim_thompson5910):

oh wait, they don't want the solutions themselves

jimthompson5910 (jim_thompson5910):

they want the number of solutions

jimthompson5910 (jim_thompson5910):

so you use D = b^2 - 4ac

OpenStudy (anonymous):

yep and i got B, wouldnt that be right?

jimthompson5910 (jim_thompson5910):

tell me what you get when you evaluate D = b^2 - 4ac

jimthompson5910 (jim_thompson5910):

plug in a = -2, b = 5, c = -3

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

D = b^2 - 4ac D = 5^2 - 4(-2)(-3) D = 1 So you were close

jimthompson5910 (jim_thompson5910):

since this discriminant is greater than 0, you get 2 real solutions

jimthompson5910 (jim_thompson5910):

so that's why the answer is C

OpenStudy (anonymous):

ohhh gotchya!